# Thermoelectric processes of quantum normal-superconductor interfaces

### v1

@@ IV.3 Chiral edge states
Recently, a peculiar thermoelectric property of edge states of a quantum Hall system (QHS) contacted to a superconducting terminal has been analyzed [237]. The electrical transport of this system is a topic of increasing interest nowadays and is the focus of many recent theoretical [238, 239, 240, 241, 242, 243, 244, 245, 246, 247, 248, 249, 250, 251, 252, 253] and experimental [254, 255, 256, 257, 258, 259, 260] works. The characteristic property is the existence of chiral Andreev states generated when the chiral edge states of the quantum Hall system get in contact to the superconductor. In the filling factor $\nu=1$, there is a single chiral edge state. When it propagates in contact to a superconductor, it becomes a chiral Andreev state, which consists in an interference of particles and holes propagating chirally.
@@
In Ref. 237, a two dimensional electron gas in the quantum Hall state with $\nu=1$ is considered in a two terminal configuration. A third terminal is defined by contacting to a superconductor with $s$-wave pairing $\Delta$ in a region of finite length. The magnetic field $B$ generating the quantum Hall effect is assumed to be perfectly screened into the superconductor. An electrical bias $eV$ is applied at one of the terminals of the quantum Hall system while the other terminal as well as the superconductor are assumed to be grounded. The following conditions are assumed: (i) $eV,k_{B}T<\Delta$ so that the transport window is fully within the superconducting gap, (ii) $eV,k_{B}T<\Delta_{B}$ so that the Zeeman splitting is larger than the energy defining the transport window. Separated electrical and thermal currents flow in response to the electrical bias: a pure electrical current flows into the superconductor, while a pure heat current flows into the grounded terminal of the QHS. Intuitively, the fact that an electrical but no heat current flows into the superconductor can be understood by noticing that under the operational conditions no quasiparticle supragap transport is expected in the superconductor (ruling out the possibility of a heat current) while, because of Andreev reflection, a charge current is expected in the superconducting terminal [152]. Interestingly, the existence of a heat current in the unbiased terminal of the QHS implies a thermoelectric response in the device. Notice that a two-terminal QHS under these operational conditions has a single pair of counter-propagating chiral edge states. This can be regarded as a perfect single-channel one-dimensional ballistic conductor with conductance $G=e^{2}/h$ and no thermoelectric response. Hence, the contact to the superconductor and the generation of a chiral Andreev state in this device effectively plays the role of a symmetry breaking mechanism enabling a nonlocal thermoelectric response. Related all-normal effects require a barrier in the conductor to generate interferences [261]. Additionally, thermal gradients in chiral edge states in the presence of superconductors may trigger intriguing $\Delta T$-noise properties [90].
@@ IV.4 Joule spectroscopy
The excess current [262] in highly transmissive NS or SNS junctions provides a fingerprint for heating and cooling mechanisms in hybrid devices  [263, 264, 265, 266]. The so-called “Joule spectroscopy” introduced in Ref. 265 is based on this property. That work considered the case of SNS junctions defined on InAl nanowires fully covered by a thin epitaxial Al layer, which exhibit a hard induced superconducting gap [267].
Figure 5: (a) Schematic of a highly transmitting SNS junction based on a hybrid superconductor-semiconductor nanowire. The temperatures of the S leads at the junction, $T_{0,i}$, can increase substantially due to the dissipated Joule power, $P_{i}$. (b) $V_{dip,i}$ are the voltages at which the S leads turn to the normal state due to Joule heating. They appear as dips in $dI/dV$, reflecting the suppression of excess current. (c) Gate dependence of $V_{dip,i}$. The dashed lines are fits to Eq. (24), confirming the $\sqrt{R_{J}}$-dependence of the dips. (d) Joule spectroscopy of a device with a floating superconducting island and two grounded superconducting leads. In this case, three $dI/dV$ dips are observed, one for each superconductor-to-normal metal transition. The magnetic field-dependence of the dips of the floating and grounded superconductors are different, owing to their distinct dominant heat dissipation mechanisms. With courtesy of Ref. 265.
The measured IV curves exhibited a series of differential conductance dips at voltages $V_{\rm dip,i}\gg 2\Delta$ [265], which could be associated with the suppression of the excess current due to the transition to the normal state of certain portions of the device (the index $i=1,2$ corresponds to these different portions). The positions of the dips were found to be tunable as a function of several control parameters such as gate voltage, magnetic field and temperature, see Fig. 5. In particular, $V_{\rm dip,i}$ were found to scale as $\sqrt{R_{J}}$, where $R_{J}$ is the normal state junction resistance, tunable by means of a gate potential acting on the junction region.
The observed features and the working principle of the technique are described by a theoretical model that relies on the balance between the heat injected on the junction by Joule effect and the different cooling mechanisms [265] (i.e. quasiparticle diffusion and electron-phonon coupling). As these mechanisms are largely inefficient at low temperatures, the junction temperature increases with respect to the bath temperature of the cryostat $T_{\rm bath}$, which might eventually lead to a transition of the leads to the normal state. For a large bias voltage $V\gg\Delta$ the junction IV curve is well described as
@@
where $T_{\rm c,i}$ is the critical temperature on lead $i$, $R_{\rm lead,i}$ is the lead normal resistance and $\Lambda\simeq 2.112$, which arises from the BCS thermal conductivity assuming $T_{\rm bath}\ll T_{\rm c,i}$. From Eqs. (22,23) and assuming a symmetric situation where $P_{1}=P_{2}\simeq V^{2}/2R_{J}$ one obtains
@@ V Two-dimensional Dirac materials
with $v_{F}$ the Fermi velocity, $\check{V}(x,y)$ the electrostatic potential, and the $2N\times 2N$ matrices $\check{\alpha}_{x,y}$ acting on the SU(2) degree of freedom defining the Dirac Hamiltonian and the $N$-dimensional space corresponding to the other degrees of freedom of the system. For example, for the surface state of a three-dimensional topological insulator (3dTI) the matrices $\check{\alpha}_{x,y}$ reduce to the Pauli matrices in spin space. By contrast, for graphene the Dirac Hamiltonian acts in sublattice space, denoted by A- and B-type atoms, but the general Hamiltonian, Eq. 26, also spans in the spin and valley degrees of freedom becoming a $8\times 8$ matrix ($N=4$).
@@
To evaluate the transport properties of hybrid junctions between Dirac materials and superconductors it is usually assumed that the contact width $W$ is much larger than the junction length $L$. Further assuming translational invariance along the contact direction, the momentum parallel to the interface between the normal and superconducting regions, $k_{y}$, is conserved, see Fig. 6. Under these approximations, the transverse momentum $k_{y}$ can be parametrized as the angle $\sin\theta=k_{y}/\mathbf{k}_{e}$, with $\mathbf{k}_{e,h}=\hbar v_{F}|E\pm E_{F}|$ the magnitude of the wavevector for electrons and holes, $v_{F}$ and $E_{F}$ respectively being the Fermi velocity and Fermi energy, and $E$ being the excitation energy. For an Andreev reflection process, the angle of the reflected hole is computed from the conservation of the transverse momentum as
@@
The charge and heat currents, Eqs. 5 and 6, include the contributions from all transverse modes. For wide junctions the sum over transverse modes results in the integral over the angle [280]
@@
with $N(E)=2W|E+E_{F}|/(hv_{F})$ the number of transverse modes. The factor $N(E)$ in Eq. 28 introduces an energy dependence unique to the two-dimensional problem involving Dirac materials. In the high-doping regime with $E_{F}\gg E$, the energy dependence of this prefactor becomes irrelevant and the angle-average resembles that of a conventional ballistic junction. By contrast, in the low-doping regime the energy dependence of the number of transport channels creates interesting asymmetries for thermoelectric effects.
@@ V.1 Graphene and graphene-like materials
Graphene is a single layer of carbon atoms arranged in a honeycomb structure, constructed as the superposition of two triangular lattices denoted by A and B atoms. As a Dirac material, low energy excitations in graphene follow a linearized Hamiltonian with two inequivalent valleys, which we assume to lie in the $x$-axis and denote them $\mathbf{K}_{\pm}=(\pm K,0)$ [279]. Around each valley, single-particle excitations follow a Dirac Hamiltonian of the form [276, 277]
@@
In the vicinity of a superconductor, a proximized graphene region presents spin and valley degeneracy, allowing us to effectively consider only one valley for electrons and the opposite for holes. For proximized structures it is also a common approximation to only consider pairing terms that are diagonal in sublattice space. Under these conditions, the DBdG Hamiltonian for graphene and graphene-like materials reduces to a $4\times 4$ matrix in sublattice and Nambu (particle-hole) spaces, namely,
@@
with the sublattice Dirac Hamiltonians defined in Eq. 30. We are considering a general pairing which could be unconventional with an anisotropic momentum dependence, but in most cases we use isotropic $s$-wave pairings $\Delta(\mathbf{k},T)=\Delta(T)$. The self-consistent temperature dependence of the parenting superconducting gap is commonly approximated by the interpolation formula $\Delta(T)\approx\Delta_{0}\tanh(1.74\sqrt{T_{c}/T-1})$, with $T_{c}$ and $\Delta_{0}$ the proximity-induced critical temperature and the zero-temperature gap amplitude, respectively. The scattering probabilities required to compute the charge and heat currents, Eqs. 5 and 6, are obtained by matching the solutions of Eq. 31 between normal ($\Delta_{0}=0$) and superconducting regions. These currents then need to be angle-averaged over all incident modes according to Eq. 28.
Early works already pointed out at the important role that specular Andreev reflections could take in transport [281]. For example, a lowly or undoped ($E_{F}\ll\Delta_{0}$) finite graphene region between two superconductors displays neutral boundary modes that propagate parallel to the normal-superconductor interfaces. These modes arise from sequential specular Andreev reflections at each interface, are sensitive to the phase difference between the superconductors, and carry energy but not charge. Consequently, the thermal conductance of such an Andreev channel becomes sensitive to the superconducting phase difference [281, 282].
Another feature of Dirac materials is the conservation of chiral symmetry in transport, which leads to Klein tunneling effects in hybrid junctions [283]. Transmission through a tunnel barrier between two Dirac materials is unusually high and oscillates with the product of the barrier height and width. This oscillatory behavior results in an unusual thermal conductance at graphene-based normal-superconductor junctions [284]. Additionally, a negative differential thermal conductance $\mathrm{d}J_{N}/\mathrm{d}(\delta T)<0$, with $\delta T$ being the thermal gradient, is also predicted for graphene NS junctions with singlet $s$-wave [285] and triplet $p$-wave superconductors [286]. The unusual transport properties of Dirac materials coupled to superconductors results in a heat current that increases as the temperature gradient decreases which can be even used to create heat current or thermal amplifiers [287]. Furthermore, electron cooling in a graphene sheet contacted by two insulator-superconductor junctions can outperform similar systems based on metallic ultra thin films [272]. Graphene’s low electron-phonon coupling preserves cooling mechanisms avoiding overheating, thus making graphene-based superconducting hybrid junctions interesting potential applications as bolometric radiation sensors [270, 271, 272].
@@ V.2 Edge states of 3D topological insulators
Another example of topological matter are Weyl semimetals, which are gapless in the bulk at discrete points and exhibit open surface states (Fermi arcs) [302]. The Weyl semimetal phase can be realized by breaking either time reversal or inversion symmetry. However, in the absence of a spin active interface Andreev reflection is suppressed due to chirality blockade for time-reversal broken Weyl semimetal/Weyl superconductor junctions. As for other Dirac materials, the thermal conductance in the low-doping regime for inversion broken junctions features an exponential dependence on the temperature and thermopower with high figure of merit around Weyl points [303]. Unusual thermal diode effects have also been predicted also for inversion symmetry-broken Weyl superconductor/Weyl semimetal Josephson junctions [304].
@@ VI Conclusion
We reviewed how heat and thermoelectrical transport are influenced by superconducting proximity effect on normal and topological materials, classifying the different phenomena on the basis of the dimensionality of the electronic degrees of freedom. In nanoscale and multiterminal zero-dimensional geometries based on quantum dots, local and nonlocal features emerge as a consequence of nonlocal correlations induced by superconducting terminals. Special attention is given to how interactions and nonlocal superconducting correlations give rise to unconventional thermoelectric phenomena. Mesoscopic one-dimensional states emerge in quantum Hall regimes, where helical modes or spin-orbit coupling can induce distinctive heat transport and thermoelectric behavior. We also reviewed how non-equilibrium Joule heating in nanowire-based Josephson junctions would affect their Josephson properties. Finally, we examined how graphene-like 2D materials, and 3D topological and Weyl materials, may exhibit unconventional responses such as thermoelectricity, anomalous thermal transport, and thermal rectification when proximized by superconductors. The discussed cases and comprehensive survey of recent literature highlight the promise of hybrid superconducting systems for thermal control, novel functionalities, and applications in superconducting quantum technologies.

### v2

@@ IV.3 Chiral edge states
Recently, a peculiar nonlocal thermoelectric property of edge states of a quantum Hall system (QHS) contacted to a superconducting terminal has been analyzed [140]. The electrical transport of this system is a topic of increasing interest nowadays and is the focus of many recent theoretical [243, 244, 245, 246, 247, 248, 249, 250, 251, 252, 253, 254, 255, 256, 257, 258] and experimental [259, 260, 261, 262, 263, 264, 265] works. The characteristic property is the existence of chiral Andreev states generated when the chiral edge states of the quantum Hall system get in contact to the superconductor. In the filling factor $\nu=1$, there is a single chiral edge state. When it propagates in contact to a superconductor, it becomes a chiral Andreev state, which consists in an interference of particles and holes propagating chirally.
@@
In Ref. 140, a two-dimensional electron gas in the quantum Hall state with $\nu=1$ is considered in a two-terminal configuration. A third terminal is defined by contacting to a superconductor with $s$-wave pairing $\Delta$ in a region of finite length. The magnetic field $B$ generating the quantum Hall effect is assumed to be perfectly screened into the superconductor. An electrical bias $eV$ is applied at one of the terminals of the quantum Hall system, while the other terminal as well as the superconductor are assumed to be grounded. The following conditions are assumed: (i) $eV,k_{B}T<\Delta$ so that the transport window is fully within the superconducting gap, (ii) $eV,k_{B}T<\Delta_{B}$ so that the Zeeman splitting is larger than the energy defining the transport window. Separated electrical and thermal currents flow in response to the electrical bias: a pure electrical current flows into the superconductor, while a pure heat current flows into the grounded terminal of the QHS. Intuitively, the fact that an electrical but no heat current flows into the superconductor can be understood by noticing that under the operational conditions no quasiparticle supragap transport is expected in the superconductor (ruling out the possibility of a heat current) while, because of Andreev reflection, a charge current is expected in the superconducting terminal [158]. Interestingly, the existence of a nonlinear heat current in the unbiased terminal of the QHS implies a nonlinear nonlocal Peltier effect in the device, despite the explicit fulfillment of the EIS in the linear regime. Notice that a two-terminal QHS under these operational conditions has a single pair of counter-propagating chiral edge states. This can be regarded as a perfect single-channel one-dimensional ballistic conductor with conductance $G=e^{2}/h$ and no thermoelectric response. Hence, the contact to the superconductor and the generation of a chiral Andreev state in this device enables the nonlocal and nonlinear thermoelectric response. Related all-normal effects require a barrier in the conductor to generate interferences [266]. Additionally, thermal gradients in chiral edge states in the presence of superconductors may trigger intriguing $\Delta T$-noise properties [93].
@@ IV.4 Joule spectroscopy
The excess current [267] in highly transmissive NS or SNS junctions provides a fingerprint for heating and cooling mechanisms in hybrid devices  [268, 269, 270, 271]. The so-called “Joule spectroscopy” introduced in Ref. 270 is based on this property. That work considered the case of SNS junctions defined on InAl nanowires fully covered by a thin epitaxial Al layer, which exhibit a hard induced superconducting gap [272].
Figure 5: (a) Schematic of a highly transmitting SNS junction based on a hybrid superconductor-semiconductor nanowire. The temperatures of the S leads at the junction, $T_{0,i}$, can increase substantially due to the dissipated Joule power, $P_{i}$. (b) $V_{dip,i}$ are the voltages at which the S leads turn to the normal state due to Joule heating. They appear as dips in $dI/dV$, reflecting the suppression of excess current. (c) Gate dependence of $V_{dip,i}$. The dashed lines are fits to Eq. (28), confirming the $\sqrt{R_{J}}$-dependence of the dips. (d) Joule spectroscopy of a device with a floating superconducting island and two grounded superconducting leads. In this case, three $dI/dV$ dips are observed, one for each superconductor-to-normal metal transition. The magnetic field-dependence of the dips of the floating and grounded superconductors are different, owing to their distinct dominant heat dissipation mechanisms. With courtesy of Ref. 270.
The measured IV curves exhibited a series of differential conductance dips at voltages $V_{\rm dip,i}\gg 2\Delta$ [270], which could be associated with the suppression of the excess current due to the transition to the normal state of certain portions of the device (the index $i=1,2$ corresponds to these different portions). The positions of the dips were found to be tunable as a function of several control parameters such as gate voltage, magnetic field and temperature, see Fig. 5. In particular, $V_{\rm dip,i}$ was found to scale as $\sqrt{R_{J}}$, where $R_{J}$ is the normal state junction resistance, tunable by means of a gate potential acting on the junction region.
The observed features and the working principle of the technique are described by a theoretical model that relies on the balance between the heat injected into the junction by the Joule effect and the different cooling mechanisms [270] (i.e. quasiparticle diffusion and electron-phonon coupling). As these mechanisms are largely inefficient at low temperatures, the junction temperature increases with respect to the bath temperature of the cryostat $T_{\rm bath}$, which might eventually lead to a transition of the leads to the normal state. For a large bias voltage $V\gg\Delta$ the junction IV curve is well described as
@@
where $T_{\rm c,i}$ is the critical temperature on lead $i$, $R_{\rm lead,i}$ is the lead normal resistance and $\Lambda\simeq 2.112$, which arises from the BCS thermal conductivity assuming $T_{\rm bath}\ll T_{\rm c,i}$. From Eqs. (26,27) and assuming a symmetric situation where $P_{1}=P_{2}\simeq V^{2}/2R_{J}$ one obtains
@@ V Two-dimensional Dirac materials
with $v_{F}$ the Fermi velocity, $\check{V}(x,y)$ the electrostatic potential, and the $2N\times 2N$ matrices $\check{\alpha}_{x,y}$ acting on the SU(2) degree of freedom defining the Dirac Hamiltonian and the $N$-dimensional space corresponding to the other degrees of freedom of the system. For example, for the surface state of a three-dimensional topological insulator (3dTI) the matrices $\check{\alpha}_{x,y}$ reduce to the Pauli matrices in spin space. By contrast, for graphene the Dirac Hamiltonian acts in sublattice space, denoted by A- and B-type atoms, but the general Hamiltonian (30) also spans in the spin and valley degrees of freedom, becoming a $8\times 8$ matrix ($N=4$).
@@
To evaluate the transport properties of hybrid junctions between Dirac materials and superconductors, it is usually assumed that the contact width $W$ is much larger than the junction length $L$. Further assuming translational invariance along the contact direction, the momentum parallel to the interface between the normal and superconducting regions, $k_{y}$, is conserved, see Fig. 6. Under these approximations, the transverse momentum $k_{y}$ can be parametrized as the angle $\sin\theta=k_{y}/\mathbf{k}_{e}$, with $\mathbf{k}_{e,h}=\hbar v_{F}|E\pm E_{F}|$ the magnitude of the wavevector for electrons and holes, $v_{F}$ and $E_{F}$ respectively being the Fermi velocity and Fermi energy, and $E$ being the excitation energy. For an Andreev reflection process, the angle of the reflected hole is computed from the conservation of the transverse momentum as
@@
The charge and heat currents, Eqs. 5 and 6, include the contributions from all transverse modes. For wide junctions, the sum over transverse modes results in the integral over the angle [285]
@@
with $N(E)=2W|E+E_{F}|/(hv_{F})$ the number of transverse modes. The factor $N(E)$ in Eq. 32 introduces an energy dependence unique to the two-dimensional problem involving Dirac materials. In the high-doping regime with $E_{F}\gg E$, the energy dependence of this prefactor becomes irrelevant and the angle-average resembles that of a conventional ballistic junction, see Sec. II.4. By contrast, in the low-doping regime, the energy dependence of the number of transport channels creates interesting asymmetries for thermoelectric effects.
@@ V.1 Graphene and graphene-like materials
Graphene is a single layer of carbon atoms arranged in a honeycomb structure, constructed as the superposition of two triangular lattices denoted by A and B atoms. As a Dirac material, low-energy excitations in graphene follow a linearized Hamiltonian with two inequivalent valleys, which we assume to lie in the $x$-axis and denote them $\mathbf{K}_{\pm}=(\pm K,0)$ [284]. Around each valley, single-particle excitations follow a Dirac Hamiltonian of the form [281, 282]
@@
In the vicinity of a superconductor, a proximized graphene region presents spin and valley degeneracy, allowing us to effectively consider only one valley for electrons and the opposite for holes. For proximized structures, it is also a common approximation to only consider pairing terms that are diagonal in sublattice space. Under these conditions, the DBdG Hamiltonian for graphene and graphene-like materials reduces to a $4\times 4$ matrix in sublattice and Nambu (particle-hole) spaces, namely,
@@
with the sublattice Dirac Hamiltonians defined in Eq. 34. We are considering a general pairing which could be unconventional with an anisotropic momentum dependence, but in most cases we use isotropic $s$-wave pairings $\Delta(\mathbf{k},T)=\Delta(T)$. The self-consistent temperature dependence of the superconducting gap is commonly approximated by the interpolation formula $\Delta(T)\approx\Delta_{0}\tanh(1.74\sqrt{T_{c}/T-1})$, with $T_{c}$ and $\Delta_{0}$ the proximity-induced critical temperature and the zero-temperature gap amplitude, respectively. The scattering probabilities required to compute the charge and heat currents, Eqs. 5 and 6, are obtained by matching the solutions of Eq. 35 between normal ($\Delta_{0}=0$) and superconducting regions. These currents then need to be angle-averaged over all incident modes according to Eq. 32.
Early works already pointed out the important role that specular Andreev reflections could take in transport [286]. For example, a lowly or undoped ($E_{F}\ll\Delta_{0}$) finite graphene region between two superconductors displays neutral boundary modes that propagate parallel to the normal-superconductor interfaces. These modes arise from sequential specular Andreev reflections at each interface, are sensitive to the phase difference between the superconductors, and carry energy but not charge. Consequently, the thermal conductance of such an Andreev channel becomes sensitive to the superconducting phase difference [286, 287].
Another feature of Dirac materials is the conservation of chiral symmetry in transport, which leads to Klein tunneling effects in hybrid junctions [288]. Transmission through a tunnel barrier between two Dirac materials is unusually high and oscillates with the product of the barrier height and width. This oscillatory behavior results in an unusual thermal conductance at graphene-based normal-superconductor junctions [289]. Additionally, a negative differential thermal conductance $\mathrm{d}J_{N}/\mathrm{d}(\delta T)<0$, with $\delta T$ being the thermal gradient, is also predicted for graphene NS junctions with singlet $s$-wave [290] and triplet $p$-wave superconductors [291]. The unusual transport properties of Dirac materials coupled to superconductors results in a heat current that increases as the temperature gradient decreases, which can even be used to create or amplify a heat current [292]. Furthermore, electron cooling in a graphene sheet contacted by two insulator-superconductor junctions can outperform similar systems based on metallic ultra-thin films [277]. Graphene’s low electron-phonon coupling preserves cooling mechanisms avoiding overheating, thus making graphene-based superconducting hybrid junctions interesting potential applications as bolometric radiation sensors [275, 276, 277].
@@ V.2 Edge states of 3D topological insulators
Another example of topological matter are Weyl semimetals, which are gapless in the bulk at discrete points and exhibit open surface states (Fermi arcs) [307]. The Weyl semimetal phase can be realized by breaking either time reversal or inversion symmetry. However, in the absence of a spin-active interface Andreev reflection is suppressed due to chirality blockade for time-reversal broken Weyl semimetal/Weyl superconductor junctions. As for other Dirac materials, the thermal conductance in the low-doping regime for inversion broken junctions features an exponential dependence on the temperature and thermopower with a high figure of merit around Weyl points [308]. Unusual thermal diode effects have also been predicted for inversion symmetry-broken Weyl superconductor/Weyl semimetal Josephson junctions [309].
@@ VI Conclusion
We reviewed how heat and thermoelectrical transport are influenced by superconducting proximity effect on normal and topological materials, classifying the different phenomena based on the dimensionality of the electronic degrees of freedom. In nanoscale and multiterminal zero-dimensional geometries based on quantum dots, local and nonlocal features emerge as a consequence of nonlocal correlations induced by superconducting terminals. Special attention is given to how interactions and nonlocal superconducting correlations give rise to unconventional thermoelectric phenomena. Mesoscopic one-dimensional states emerge in quantum Hall regimes, where helical modes or spin-orbit coupling can induce distinctive heat transport and thermoelectric behavior. We also reviewed how non-equilibrium Joule heating in nanowire-based Josephson junctions would affect their Josephson properties. Finally, we examined how graphene-like 2D materials, and 3D topological and Weyl materials, may exhibit unconventional responses such as thermoelectricity, anomalous thermal transport, and thermal rectification when proximized by superconductors.
The discussed cases and comprehensive survey of recent literature highlight the promise of hybrid superconducting systems for thermal control, novel functionalities, and applications in superconducting quantum technologies, taking the form of power generators or refrigerators that exploit or mitigate undesired hotspots, or of thermal diodes, transistors or switches able to operate with millikelvin temperature differences through nanoscale distances.